OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line (8,-4);6x-5y=8 the equation of the line is y= simplify your answer in the form y=mx+b using integers or fractions

OpenStudy (aroub):

5y=6x+8 y=6/5x+8/5 y=mx+b (8,-4) -4=6/5(8)+b -4=48/5+b -20/5-48/5=b -68/5 =b y=6/5x-68/5

myininaya (myininaya):

6x-5y=8 6x-8=5y 6/5*x-8/5=y slope is 6/5 so slope of \\ line is 6/5 y=6/5*x+b now we need y-intercept(aka b) we know a point on the line (x=8,y=-4) -4=6/5*8+b -4=48/5+b -4-48/5=b -20/5-48/5=b -68/5=b so the line we are looking for is y=6/5*x-68/5

OpenStudy (anonymous):

so the answer is y=6/5x-68/5?

OpenStudy (aroub):

can u give a good answer please , thank you! :)

OpenStudy (anonymous):

i was working on it thanx

OpenStudy (aroub):

thank you very much! :)